OpenStudy (anonymous):

do anyone know how to do this 729.8 17.8+6.4 2

OpenStudy (anonymous):

like i think i know the formula A l+w =h but u have to solve for l 2

OpenStudy (anonymous):

find L if A= 729.8in , h=17.8, and w=6.4in

OpenStudy (shane_b):

Is that A supposed to be Area or Volume? I ask because you have a length, width and a height shown there. If it's asking you to find the length and you have the area and width, you can simply use A = L * W.

OpenStudy (anonymous):

yes

OpenStudy (shane_b):

\[A = L * W\]\[729.8in^2 = L * 6.4in\]\[\frac{729.8in^2}{6.4in}=L\]I don't see why you were given a height if this is what is being asked.

OpenStudy (anonymous):

but it said that the h= 17.8

OpenStudy (shane_b):

You're given the area and the width...you don't need the height at all to find the length.

OpenStudy (anonymous):

ooo....

OpenStudy (anonymous):

thanks